OpenStudy (anonymous):

1. What is the area of a trapezoid with height 5 m and bases 8 m and 1 m? 6.5 m² 22.5 m² 24 m² 45 m²

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

no, the trapezoid formula is 1/2 * 5 * ( 8 +1)

OpenStudy (anonymous):

22.5 - B

OpenStudy (anonymous):

i think it is \(\frac{1}{2}(b_1+b_2)\times h\)

OpenStudy (anonymous):

and ur correct @satellite73, but wat else is new :P

OpenStudy (anonymous):

\[\frac{1}{2}(8+1)\times 5\]

OpenStudy (anonymous):

ok so it is B sorry i was testing yesterday

OpenStudy (anonymous):

yes, B is the answer